I am not able to understand the shape of orbitals and the concept of hybridisation. Please explain it to me in the best possible way you can

Grade:

1 Answers

vikas askiitian expert
13 years ago

to find hybridisation of any molecule ,ion use HON rule

according to this

                     x = (number of electron in valence shell of central atom + no of monovalent atom - charge)/2

if X = 2 , hyberdisation is SP

if X = 3 , hyberdisation is SP2 & so on

eg) NH4+

   X =  (5+ 4 - 1)/2 = 4  = SP3

H2O

   X = (6+2)/2=4 = SP3

CO2

  X=(4)/2 = 2 = SP